# Casper: Causality-Aware Spatiotemporal Graph Neural Networks for Spatiotemporal Time Series Imputation

### v1

@@ Abstract.
Spatiotemporal time series is the foundation of understanding human activities and their impacts, which is usually collected via monitoring sensors placed at different locations. The collected data usually contains missing values due to various failures, e.g., mechanical damage, which have significant impact on data analysis. To impute the missing values, a lot of methods have been introduced in the literature. When recovering a specific data point, most existing methods tend to take into consideration all the information relevant to that point regardless of whether they have a cause-and-effect relationship. During data collection, it is inevitable that some unknown confounders are included, e.g., background noise in time series and non-causal shortcut edges in the constructed sensor network. These confounders could open backdoor paths between the input and output – in other words, they establish non-causal correlations between the input and output. Over-exploiting these non-causal correlations could result in overfitting and make the model vulnerable to noises. In this paper, we first revisit spatiotemporal time series imputation from a causal perspective, which shows the causal relationships among the input, output, embeddings and confounders. Next, we show how to block the confounders via the frontdoor adjustment. Based on the results of the frontdoor adjustment, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which contains a novel Spatiotemporal Causal Attention (SCA) and a Prompt Based Decoder (PBD). PBD could reduce the impact of confounders and SCA could discover the sparse causal relationships among embeddings. Theoretical analysis reveals that SCA discovers causal relationships based on the values of gradients. We evaluate Casper on three real-world datasets, and the experimental results show that Casper could outperforms the baselines and could effectively discover the causal relationships.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Spatiotemporal time series is the cornerstone of analyzing and understanding the patterns and impact of human activities, e.g., traffic flows (Li et al., 2018) and greenhouse gas emissions (Zheng et al., 2015), which is the basis for formulating public policies, e.g., the traffic transportation policy and the clean air act. Spatiotemporal time series is usually collected by monitoring sensors. Traffic flow time series, e.g., speed and occupancy, is recorded by the radar sensors placed along roads. Air pollution time series, e.g., concentrations of CO and PM2.5, is collected from air quality monitoring sites across cites.
@@
In the real world, it is not uncommon that the collected spatiotemporal time series is incomplete with a lot of missing data due to various failures, e.g., the Internet is disconnected or the sensors have mechanical damages. The missing data usually has significant impact on the process and conclusion of data analysis. Therefore, how to reconstruct the missing data from the observed data, i.e., imputation, is a fundamental problem of spatiotemporal time series analysis. In recent years, deep learning methods become the mainstream for time series imputation. Most existing deep time series imputation methods (Cao et al., 2018; Luo et al., 2018, 2019) use Recurrent Neural Network (RNN) to capture the temporal dynamics of time series and autoregressively recover the missing data by the predicted values. Recent deep learning methods (Yıldız et al., 2022) propose to use non-autoregressive structures, e.g., Transformer (Vaswani et al., 2017), to avoid the progressive error propagation incurred via the autoregression in RNN by concurrently considering the entire input context. However, these methods only consider the temporal patterns yet overlook the spatial relationships among sensors, e.g., geographical distances. To further account for spatial relationships, spatiotemporal graph neural networks are introduced (Jing et al., 2021; Huang et al., 2020; Marisca et al., 2022). Although these methods have achieved impressive performance for recovering the missing values, when recovering a data point, they tend to include all the available information related to the data point as reference without distinguishing whether there is a causal relationship between them.
@@
When collecting datasets, it is inevitable that some unknown confounders (Pearl and Mackenzie, 2018) are included. For example, the background noise might be recorded, and non-causal shortcut edges might be built for two sensors. Let’s take the air monitoring sensor network as a concrete example to understand what are non-causal edges. A common practice to build the network is adding an edge for two sensors if their distance is below a threshold (Marisca et al., 2022). Although simple and usually effective, the distance based network does not necessarily imply the real causality between sensors. In the real world, air flow between two locations could be influenced by other factors, e.g., wind direction and terrain. An example is shown in Figure 6 in the Experiments section. Simply exploiting the shortcut edges without discovering the causality, could make the model overfit the training data and be vulnerable to the noise during the test.
@@
To reduce the negative effects brought by the confounders, we first review the process of spatiotemporal time series imputation from a causal perspective and show the causal relationships among the input, output, embeddings and confounders. As shown in Figure 1, the confounders could open shortcut backdoor paths between the input and output. Then, we show how to eliminate the backdoor paths opened by confounders via the frontdoor adjustment (Pearl and Mackenzie, 2018). Based on the results of the frontdoor adjustment, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which is equipped with a novel Spatiotemporal Causal Attention (SCA) and a Prompt Based Decoder (PBD). The proposed PBD effectively reduces the impact of the unknown confounders by injecting the global context information of datasets to the embeddings. Inspired by (Jia et al., 2022), which uses learnable prompts to capture the contextual information of downstream tasks when tuning visual models, PBD leverages prompts to learn the contextual information of datasets automatically rather than employing external models to approximate the context. To further enforce sparse causality between embeddings, we introduce SCA, which determines the cause-and-effect relationship via a causal gate. It can be theoretically proven that the proposed causal gate (1) enforces the sparsity since it converges to 0 or 1; (2) is actually a gradient-based explanation similar to (Selvaraju et al., 2017), which determines the causality based on the values of gradients. We extensively evaluate Casper on three real-world spatiotemporal time series imputation datasets. The experimental results show that Casper could significantly outperform baselines and could effectively discover the causality.
The major contributions are summarized as follows:
@@
  • We propose a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper) based on the frontdoor adjustment. Casper is equipped with a novel Spatiotemporal Causal Attention (SCA) and a Prompt Based Decoder (PBD). PBD effectively blocks the backdoor paths and SCA explicitly reveals the causality between emebeddings.
@@ 2. Preliminary
Let $\mathbf{X}\in\mathbb{R}^{N\times T}$ be the values of past $T$ steps of $N$ time series, and $\hat{y}_{i,T+1}=f_{i}(\mathbf{X})\in\mathbb{R}$ be the time series forecasting function predicting the future value of the $i$-th time series at the $T+1$ step, where $i\in[1,\dots,N]$. The $j$-th time series is said to Granger cause the $i$-th time series if there exists a point $x^{\prime}_{j,t}\neq x_{j,t}$, $t\in[1,\dots,T]$, such that $f_{i}(\mathbf{X}^{\prime})\neq f_{i}(\mathbf{X})$, where $\mathbf{X}^{\prime}$ is obtained by replacing $x_{j,t}$ in $\mathbf{X}$ with $x^{\prime}_{j,t}$.
Generally speaking, if $x_{j,t}$ has impact on the prediction of the future value of the $i$-th time series, then the $j$-th time series Granger causes the $i$-th time series. In the case that $f_{i}$ is a linear model:
(1) $\hat{y}_{i,T+1}=f_{i}(\mathbf{X})=\sum_{i^{\prime},t^{\prime}}\alpha_{i^{\prime},t^{\prime}}x_{i^{\prime},t^{\prime}}.$
If the coefficient $\alpha_{j,t}\neq 0$, then the $j$-th time series Granger causes the $i$-th time series.
@@ 3. Methodology
In this section, we first provide a causal view of the spatiotemporal imputation task, and show how to eliminate the unknown confounds by frontdoor adjustment, based on which, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper). Finally, we provide further analysis of Casper.
@@ 3.1. Causal View of Spatiotemporal Imputation
Given an incomplete spatiotemporal time series $\mathbf{G}=(\mathbf{X},\mathbf{A},\mathbf{M})$, a standard deep imputation model $f=f^{D}\circ f^{E}$, where $f^{E}$, $f^{D}$ are the encoder and decoder, works as follows: 1) $f^{E}$ extracts embeddings =$\mathcal{H}=\{\mathbf{h}_{i,t}\}_{i=1,t=1}^{N,T}$ from $\mathbf{G}$, 2) $f^{D}$ generates predictions $\{\hat{y}_{i,t}\}_{i=1,t=1}^{N,T}$ based on $\mathcal{H}$ to recover $\mathcal{Y}=\{{y}_{i,t}\}_{i=1,t=1}^{N,T}$. The model $f$ is trained by a reconstruction error e.g., MAE or RMSE. Since minimizing MAE (or RMSE) is equivalent to maximizing the log-likelihood of Laplace (or Gaussian) distribution (Hodson, 2022), and thus we can view the objective of spatiotemporal imputation as maximizing $P(\mathcal{Y}|\mathcal{G})$. Most existing studies focus on maximizing $P(\mathcal{Y}|\mathcal{G})$ yet few discuss the cause-and-effect relationship between $\mathcal{G}$ and $\mathcal{Y}$. In this paper, we study their causality based on the Structure Causal Model (SCM) (Pearl and Mackenzie, 2018).
@@
3.1.1. Structure Causal Model
During data collection, it is inevitable that some unknown confounders $\mathcal{C}$ are included in the datasets, which influence both $\mathcal{G}$ and $\mathcal{Y}$. For example, sensors might record random background noise, and the constructed sensor network might contain shortcut edges. The undesired information might bridge the input $\mathcal{G}_{M}$ and the output $\mathcal{Y}$ with spurious correlations, which could lead to overfitting and make the model error-prone.
@@ 3.1.1. Structure Causal Model
3.1.2. Frontdoor Adjustment
In statistics (Murphy, 2012), a simple way to exclude the variable $\mathcal{C}$ in the SCM in Figure 1 is to marginalize it out. However, marginalization requires $\mathcal{C}$ to be observable and measured by the marginal distribution $P(\mathcal{C})$, but in spatiotemporal imputation, $\mathcal{C}$ is usually unknown and difficult to measure. Rather than directly marginalizing $\mathcal{C}$ out, we resort to the frontdoor adjustment (Pearl and Mackenzie, 2018), which uses Pearl’s do-calculus (Pearl and Mackenzie, 2018) to block the backdoor paths. We follow the three steps of the frontdoor adjustment.
@@ 3.1.2. Frontdoor Adjustment
In general, $P(\mathbf{h}_{i,t}|\mathcal{G})$ can be viewed as the encoder $f^{E}$, and the rest part, including $\sum_{\mathbf{G}}P(\mathcal{Y}|\mathbf{h}_{i,t},\mathcal{G})P(\mathbf{G})$ and the sum over all data points $\sum_{i,t}$, can be viewed as the decoder $f^{D}$. In the next subsection, we show how to implemet Equation (5).
@@ 3.2. Architecture of Casper
In this subsection, based on Equation (5), we propose a novel Cusality-Aware Spatiotemporal Graph Neural Network (Casper). We first present an overview of Casper  which comprised of a Prompt Based Decoder (PBD) and an encoder with Spatiotemporal Causal Attention (SCA). Next, we elaborate PBD and SCA in detail.
@@ 3.2.1. Overview
The final embeddings of the $l$-th layer is given by:
@@
3.2.2. Prompt Based Decoder
Suppose we are given the input $\mathbf{G}$ and the target is to recover $\mathcal{Y}=y_{i,t}$. In Equation (5), the decoder $f^{D}$ is comprised of (1) a sum over all possible $\mathbf{G}^{\prime}$ in the dataset $\sum_{\mathbf{G}^{\prime}}P(y_{i,t}|\mathbf{h}_{i^{\prime},t^{\prime}},\mathbf{G}^{\prime})P(\mathbf{G}^{\prime})$ and (2) a sum over all data points $\sum_{i^{\prime},t^{\prime}}$ in $\mathbf{G}$. For (2), since the encoders nowadays, e.g., Transformer (Vaswani et al., 2017), are so powerful that could encode sufficient context information $\mathbf{G}$ in $\mathbf{h}_{i,t}$, and thus the decoder $f^{D}$ could only take $\mathbf{h}_{i,t}$ as input (Marisca et al., 2022), instead of all possible $\mathbf{h}_{i^{\prime},t^{\prime}}$. Therefore, we could drop $\sum_{i^{\prime},^{\prime}t}$ and only implement $\sum_{\mathbf{G}^{\prime}}P(y_{i,t}|\mathbf{h}_{i,t},\mathbf{G}^{\prime})P(\mathbf{G}^{\prime})$.
@@ 3.2.2. Prompt Based Decoder
An illustration of the proposed PBD is shown in Figure 3. $\mathcal{Q}=\mathbf{h}_{i,t}$ is the query, $\mathcal{P}=\{\mathbf{p}_{n}\}_{n=1}^{N_{P}}$ is the set of learnable prompts. In Figure 3, the Project is a linear function followed with a LayerNorm (Ba et al., 2016). Details of the scaled dot-product attention can be found in (Vaswani et al., 2017), and it can be easily extended into multi-head version as in (Vaswani et al., 2017).
3.2.3. Spatiotemporal Causal Attention
Attention functions (Definition 2.4) have become indispensable in deep learning models (Vaswani et al., 2017), which could effectively capture the context information for the target embedding. Although attention scores could show the correlation between embeddings, yet correlation does not necessarily imply causality and thus sometimes could induce undesired non-causal information into embeddings (Sui et al., 2022). Based on the frontdoor adjustment, PBD could eliminate the impact of the unknown confounders $\mathcal{C}$ and ensure the causality between the input $\mathcal{G}$ and the output $\mathcal{Y}$ by summing over $\mathbf{G}$, $i$ and $t$. However, it guides the attention functions to discover the causal relationships at a high level, and thus the learned causal relationships, i.e., attention scores, might still be dense and a little bit difficult to interpret. To directly guide the model to discover the sparse causal relationships, we first define the causality for embeddings (Definition 3.1) based on the Granger causality (Granger, 1969) (Definition 2.3), and then introduce a novel Spatiotemporal Causal Attention (SCA) module to discover the sparse causality between embeddings.
Definition 0 (Granger Causality for Embeddings).
Denote the target embedding and the set of context embeddings as $\mathcal{Q}=\mathbf{h}_{i,t}^{in}$ and $\mathcal{K}=\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\}$. Let $\mathbf{h}_{i,t}^{out}=f_{i,t}(\mathbf{h}_{i,t}^{in};\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\})$ be an embedding updating function. If there is a $\mathbf{h}_{j,r}^{in}\in\mathcal{K}$ s.t. changing the value of $\mathbf{h}_{j,r}^{in}$ will change the value of $\mathbf{h}_{i,t}^{out}$, then $\mathbf{h}_{j,r}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$.
@@ 3.2.3. Spatiotemporal Causal Attention
(12) $\mathbf{h}_{i,t}^{out}=\sum_{i^{\prime},t^{\prime}}\alpha_{i^{\prime},t^{\prime}}\mathbf{v}_{i^{\prime},t^{\prime}},\quad\mathbf{v}_{i^{\prime},t^{\prime}}=f_{v}(\mathbf{h}_{i,t}^{in},\mathbf{h}_{i^{\prime}t^{\prime}}^{in}),$
where $\alpha_{i^{\prime},t^{\prime}}$ is the attention weight, and $f_{v}$ is the message function. According to Definition 3.1, if $\alpha_{j,r}\neq 0$, then $\mathbf{h}_{j,r}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$; otherwise if $\alpha_{j,r}=0$, then $\mathbf{h}_{j,r}^{in}$ does not Granger cause $\mathbf{h}_{i,t}^{out}$. In practice, without directly manipulating, $\alpha_{i^{\prime},t^{\prime}}>0$ holds for many noise messages $\mathbf{v}_{i^{\prime},t^{\prime}}$, as shown in Figure 4(c)-4(d) in our experiments.
@@
Let $\mathcal{Q}=\mathbf{h}_{i,t}^{in}$ be the target query embedding, and $\mathcal{N}(i)$ be the neighbors of the $i$-th sensor, i.e., $\mathbf{A}[i,i^{\prime}]\neq 0$. Denote the context keys as $\mathcal{K}=\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\}$ and the message values from $\mathcal{N}(i)$ to the point $(i,t)$ as $\mathcal{V}=\{\mathbf{v}_{i^{\prime},t^{\prime}}\}$, where $i^{\prime}\in\mathcal{N}(i)$, $t^{\prime}\in[1,\cdots,T]$ and $\mathbf{v}_{i^{\prime},t^{\prime}}=\text{MLP}([\mathbf{h}_{i,t}^{in};\mathbf{h}_{i^{\prime},t^{\prime}}^{in}])$. We define SCA as:

### v3

# Causality-Aware Spatiotemporal Graph Neural Networks for Spatiotemporal Time Series Imputation

@@ Abstract.
Spatiotemporal time series are usually collected via monitoring sensors placed at different locations, which usually contain missing values due to various failures, such as mechanical damages and Internet outages. Imputing the missing values is crucial for analyzing time series. When recovering a specific data point, most existing methods consider all the information relevant to that point regardless of the cause-and-effect relationship. During data collection, it is inevitable that some unknown confounders are included, e.g., background noise in time series and non-causal shortcut edges in the constructed sensor network. These confounders could open backdoor paths and establish non-causal correlations between the input and output. Over-exploiting these non-causal correlations could cause overfitting. In this paper, we first revisit spatiotemporal time series imputation from a causal perspective and show how to block the confounders via the frontdoor adjustment. Based on the results of frontdoor adjustment, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which contains a novel Prompt Based Decoder (PBD) and a Spatiotemporal Causal Attention (SCA). PBD could reduce the impact of confounders and SCA could discover the sparse causal relationships among embeddings. Theoretical analysis reveals that SCA discovers causal relationships based on the values of gradients. We evaluate Casper on three real-world datasets, and the experimental results show that Casper could outperform the baselines and could effectively discover the causal relationships.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679642isbn: 979-8-4007-0436-9/24/10ccs: Information systems Data miningccs: Information systems Sensor networks
@@ 1. Introduction
Spatiotemporal data mining (Atluri et al., 2018) is the cornerstone of analyzing and understanding the patterns of spacetime and human activities, such as environmental monitoring (Li et al., 2018; Roach et al., 2020; Zheng et al., 2015; Marisca et al., 2022), e-business (Kang and McAuley, 2018; Zhou et al., 2020; Jing et al., 2024b, [n. d.]; Yan et al., 2024c; Li et al., 2022; Jing et al., 2021a; Yan et al., 2022) and social science (Zhao and Tang, 2017; Du et al., 2021; Yan et al., 2021b, 2024b; Zeng et al., 2023, 2024a; Feng et al., 2022, 2024; Zhou, 2023). Time series (Jing et al., 2024a; Wang et al., 2023; Li et al., 2021; Zheng et al., 2024a) is one of the most common data types, which is usually collected by monitoring sensors. For example, traffic flow time series (Li et al., 2018), e.g., speed, is recorded by the radar sensors on roads. Air pollution time series (Zheng et al., 2015), e.g., concentrations of PM2.5, is collected from air quality monitoring sites across cities.
@@
In the real world, it is not uncommon that the collected spatiotemporal time series is incomplete with missing data due to various failures, e.g., sensors have mechanical damage. The missing data usually significantly impacts the process and conclusion of data analysis. Therefore, how to reconstruct the missing data from the observed data, i.e., imputation, is a fundamental problem of spatiotemporal time series analysis. In recent years, deep learning methods become the mainstream for time series imputation. Most existing deep time series imputation methods (Cao et al., 2018; Luo et al., 2018, 2019) use Recurrent Neural Network (RNN) to capture the temporal dynamics of time series and autoregressively recover the missing data by the predicted values. Recent deep learning methods (Yıldız et al., 2022) propose to use non-autoregressive structures, e.g., Transformer (Vaswani et al., 2017), to avoid the progressive error propagation incurred via the autoregression in RNN by concurrently considering the entire input context. However, these methods only consider the temporal patterns yet overlook the spatial relationships among sensors, e.g., geographical distances. To further account for spatial relationships, graph neural networks (Kipf and Welling, 2016; Veličković et al., 2017; Yan et al., 2024a; Zeng et al., 2024b; Xu et al., 2024; Zheng et al., 2021) are extended to the spatiotemporal setting (Jing et al., 2021b; Huang et al., 2020; Marisca et al., 2022; Wang et al., 2023). Although these methods have achieved impressive performance in recovering the missing values, they tend to include all the available information related to the missing point as references without distinguishing whether there is a causal relationship between them.
@@
When collecting datasets, it is inevitable to include some unknown confounders (Pearl and Mackenzie, 2018). For example, the background noise might be recorded, and non-causal shortcut edges might be built for two sensors. Let’s take the air monitoring sensor network as a concrete example to understand the non-causal edges. A common practice to build the network is adding an edge for two sensors if their distance is below a threshold (Marisca et al., 2022). Although simple and usually effective, the distance-based network does not necessarily imply the real causality between sensors. In the real world, air flow between two locations could be influenced by other factors, e.g., wind direction and terrain. An example is shown in Figure 6 in Section 4.4. Simply exploiting the shortcut edges without discovering the causality could make the model overfit the training data and be vulnerable to noise during inference.
@@
To reduce the negative effects brought by confounders, we first review the process of spatiotemporal time series imputation from a causal perspective (Pearl and Mackenzie, 2018) to show the causal relationships among the input, output, embeddings, and confounders. The results show that confounders could establish undesired non-causal shortcut backdoor paths between the input and output. Then, we show how to eliminate the backdoor paths via the frontdoor adjustment (Pearl and Mackenzie, 2018). Based on the results of the frontdoor adjustment, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which is equipped with a novel Prompt Based Decoder (PBD) and a Spatiotemporal Causal Attention (SCA). The proposed PBD effectively reduces the impact of unknown confounders by injecting the global context information of datasets into the embeddings. Inspired by (Jia et al., 2022), which uses learnable prompts to capture the contextual information of downstream tasks when tuning visual models, PBD leverages prompts to learn the contextual information of datasets automatically rather than employing external models to approximate the context. To further enforce sparse causality between embeddings, we introduce SCA, which determines the cause-and-effect relationship via a causal gate. It can be theoretically proven that the proposed causal gate (1) enforces the sparsity since it converges to 0 or 1; (2) is a gradient-based explanation similar to (Selvaraju et al., 2017), which determines the causality based on the values of gradients. We extensively evaluate Casper on three real-world datasets. The experimental results show that Casper could significantly outperform baselines and could effectively discover causality.
The major contributions of the paper are summarized as follows:
@@
  • We propose a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper) based on the frontdoor adjustment. Casper is equipped with a novel Spatiotemporal Causal Attention (SCA) and a Prompt Based Decoder (PBD). PBD effectively blocks the backdoor paths and SCA explicitly reveals the causality between embeddings.
@@ 2. Preliminary
Let $\mathbf{X}\in\mathbb{R}^{N\times T}$ be the values of past $T$ steps of $N$ time series, and $\hat{y}_{i,T+1}=f_{i}(\mathbf{X})\in\mathbb{R}$ be the time series forecasting function predicting the future value of the $i$-th time series at the $T+1$ step, where $i\in\{1,\dots,N\}$. The $i^{\prime}$-th time series is said to Granger cause the $i$-th time series if there exists a point $x^{\prime}_{i^{\prime},t^{\prime}}\neq x_{i^{\prime},t^{\prime}}$, $t^{\prime}\in\{1,\dots,T\}$, such that $f_{i}(\mathbf{X}^{\prime})\neq f_{i}(\mathbf{X})$, where $\mathbf{X}^{\prime}$ is obtained by replacing $x_{i^{\prime},t^{\prime}}$ in $\mathbf{X}$ with $x^{\prime}_{i^{\prime},t^{\prime}}$.
Generally, if $x_{i^{\prime},t^{\prime}}$ impacts the prediction of the future value of the $i$-th time series, then the $i^{\prime}$-th time series Granger causes the $i$-th time series. In the case that $f_{i}$ is a linear model:
(1) $\hat{y}_{i,T+1}=f_{i}(\mathbf{X})=\sum_{i^{\prime}=1,t^{\prime}=1}^{N,T}\alpha_{i^{\prime},t^{\prime}}x_{i^{\prime},t^{\prime}}.$
if the coefficient $\alpha_{i^{\prime},t^{\prime}}\neq 0$, then the $i^{\prime}$-th time series Granger causes the $i$-th time series.
@@ 3. Methodology
In this section, we first provide a causal view of the spatiotemporal imputation task, and show how to eliminate the impact of unknown confounds by frontdoor adjustment, based on which, we introduce a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper). Finally, we provide further analysis of Casper.
@@ 3.1. Causal View of Spatiotemporal Imputation
Given an incomplete spatiotemporal time series $\mathbf{G}=(\mathbf{X},\mathbf{A},\mathbf{M})$, a standard deep imputation model $f=f^{D}\circ f^{E}$, where $f^{E}$, $f^{D}$ are the encoder and decoder, works as follows: (1) $f^{E}$ extracts embeddings =$\mathcal{H}=\{\mathbf{h}_{i,t}\}_{i=1,t=1}^{N,T}$ from $\mathbf{G}$, (2) $f^{D}$ generates predictions $\{\hat{y}_{i,t}\}_{i=1,t=1}^{N,T}$ based on $\mathcal{H}$ to recover $\mathcal{Y}=\{{y}_{i,t}\}_{i=1,t=1}^{N,T}$. The model $f$ is trained by a reconstruction error e.g., MAE or RMSE. Since minimizing MAE (or RMSE) is equivalent to maximizing the log-likelihood of Laplace (or Gaussian) distribution (Hodson, 2022), and thus we can view the objective of spatiotemporal imputation as maximizing $P(\mathcal{Y}|\mathcal{G})$. Most existing studies focus on maximizing $P(\mathcal{Y}|\mathcal{G})$ yet few discuss the cause-and-effect relationship between $\mathcal{G}$ and $\mathcal{Y}$. In this paper, we study their causality based on the Structure Causal Model (SCM) (Pearl and Mackenzie, 2018).
@@
Structure Causal Model
During data collection, it is inevitable that some unknown confounders $\mathcal{C}$ are included in datasets, which influence both $\mathcal{G}$ and $\mathcal{Y}$. For example, sensors might record random background noise, and the constructed sensor network might contain shortcut edges. The undesired information might bridge the input $\mathcal{G}_{M}$ and the output $\mathcal{Y}$ with spurious correlations, which could lead to overfitting and make the model error-prone.
@@
Frontdoor Adjustment
In statistics (Murphy, 2012), a simple way to exclude the variable $\mathcal{C}$ in the SCM in Figure 1 is to marginalize it out. However, marginalization requires $\mathcal{C}$ to be observable and measured by the marginal distribution $P(\mathcal{C})$, but in spatiotemporal imputation, $\mathcal{C}$ is usually unknown and difficult to measure. Rather than directly marginalizing $\mathcal{C}$ out, we resort to the frontdoor adjustment (Pearl and Mackenzie, 2018), which uses Pearl’s do-calculus (Pearl and Mackenzie, 2018) to block the backdoor paths. We follow the three steps of the frontdoor adjustment as follows.
@@
In general, $P(\mathbf{h}_{i,t}|\mathcal{G})$ can be viewed as the encoder $f^{E}$, and the rest part, including $\sum_{\mathbf{G}}P(\mathcal{Y}|\mathbf{h}_{i,t},\mathcal{G})P(\mathbf{G})$ and the sum over all data points $\sum_{i,t}$, can be viewed as the decoder $f^{D}$. In the next subsection, we show how to implement Equation (5).
@@ 3.2. Architecture of Casper
In this subsection, based on Equation (5), we propose a novel Cusality-Aware Spatiotemporal Graph Neural Network (Casper). We first present an overview of Casper, consisting of a Prompt Based Decoder (PBD) and an encoder with Spatiotemporal Causal Attention (SCA). Next, we elaborate PBD and SCA in detail.
@@
The final embeddings of the $l$-th layer are given by:
@@
Prompt Based Decoder
Suppose we are given the input $\mathbf{G}$ and the target is to recover $\mathcal{Y}=y_{i,t}$. In Equation (5), the decoder $f^{D}$ is comprised of (1) a sum over all possible $\mathbf{G}^{\prime}$ in the dataset $\sum_{\mathbf{G}^{\prime}}P(y_{i,t}|\mathbf{h}_{i^{\prime},t^{\prime}},\mathbf{G}^{\prime})P(\mathbf{G}^{\prime})$ and (2) a sum over all data points $\sum_{i^{\prime},t^{\prime}}$ in $\mathbf{G}$. For (2), since the encoders nowadays, e.g., Transformer (Vaswani et al., 2017), are so powerful that could encode sufficient context information $\mathbf{G}$ in $\mathbf{h}_{i,t}$, and thus the decoder $f^{D}$ could only take $\mathbf{h}_{i,t}$ as input (Marisca et al., 2022), instead of all possible $\mathbf{h}_{i^{\prime},t^{\prime}}$. Therefore, we could drop $\sum_{i^{\prime},t^{\prime}}$ and only implement $\sum_{\mathbf{G}^{\prime}}P(y_{i,t}|\mathbf{h}_{i,t},\mathbf{G}^{\prime})P(\mathbf{G}^{\prime})$.
@@
An illustration of the proposed PBD is shown in Figure 3. $\mathcal{Q}=\mathbf{h}_{i,t}$ is the query, $\mathcal{P}=\{\mathbf{p}_{n}\}_{n=1}^{N_{P}}$ is the set of learnable prompts, which are randomly initialized embedding vectors. In Figure 3, the Project is a linear function followed with a LayerNorm (Ba et al., 2016). Details of the scaled dot-product attention can be found in (Vaswani et al., 2017), and it can be easily extended into the multi-head version as in (Vaswani et al., 2017).
Spatiotemporal Causal Attention
Attention functions (Definition 2.4) have become indispensable in deep learning models (Vaswani et al., 2017; Jing et al., 2022a), which could effectively capture the context information for the target embedding. Although attention scores could show the correlation between embeddings, correlation does not necessarily imply causality and thus sometimes could induce undesired non-causal information into embeddings (Sui et al., 2022). Based on the frontdoor adjustment, PBD could eliminate the impact of unknown confounders $\mathcal{C}$ and ensure the causality between the input $\mathcal{G}$ and the output $\mathcal{Y}$ by summing over $\mathbf{G}$, $i$ and $t$. However, it guides the attention functions to discover the causal relationships at a high level, and thus the learned causal relationships, i.e., attention scores, might still be dense and a little bit difficult to interpret (see Figuers 5(c) and 5(g)). To directly guide the model to discover the sparse causal relationships, we first define the causality for embeddings (Definition 3.1) based on the Granger causality (Granger, 1969) (Definition 2.3), and then introduce a novel Spatiotemporal Causal Attention (SCA) module to discover the sparse causality between embeddings.
Definition 0 (Unconstrained Granger Causality for Embeddings).
Denote the target embedding as $\mathcal{Q}=\mathbf{h}_{i,t}^{in}$ and the set of context embeddings as $\mathcal{K}=\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\}_{i^{\prime}=1,t^{\prime}=1}^{N,T}$. Let $\mathbf{h}_{i,t}^{out}=f_{i,t}(\mathcal{Q};\mathcal{K})$ be an embedding updating function, e.g., attention function. If there is a $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\in\mathcal{K}$ s.t. changing the value of $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ will change the value of $\mathbf{h}_{i,t}^{out}$, then $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$.
We do not strictly enforce the time $t^{\prime}\leq t$ for the imputation task as for the forecasting task (Definition 2.3), since (1) a missing value could appear at the beginning of the input time series segment, and there are no prior points available; (2) most imputation methods in the literature consider both past $t^{\prime}\leq t$ and future $t^{\prime}>t$ reference points $\mathbf{h}_{i^{\prime},t^{\prime}}$ for the data point to be imputed $\mathbf{h}_{i,t}$; (3) given the learned weight $w_{i^{\prime},t^{\prime}}$ between $\mathbf{h}_{i^{\prime},t^{\prime}}$ and $\mathbf{h}_{i,t}$, it is easy to distinguish whether it is from the past or future by comparing $t^{\prime}$ and $t$, and thus we can easily obtain the time-constrained causal graph if necessary.
@@
(12) $\mathbf{h}_{i,t}^{out}=\sum_{i^{\prime}=1,t^{\prime}=1}^{N,T}\alpha_{i^{\prime},t^{\prime}}\mathbf{v}_{i^{\prime},t^{\prime}},\quad\mathbf{v}_{i^{\prime},t^{\prime}}=f_{v}(\mathbf{h}_{i,t}^{in},\mathbf{h}_{i^{\prime}t^{\prime}}^{in}),$
where $\alpha_{i^{\prime},t^{\prime}}$ is the attention weight, and $f_{v}$ is the message function. According to Definition 3.1, if $\alpha_{i^{\prime},t^{\prime}}\neq 0$, then $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ Granger causes $\mathbf{h}_{i,t}^{out}$; otherwise $\mathbf{h}_{i^{\prime},t^{\prime}}^{in}$ does not Granger cause $\mathbf{h}_{i,t}^{out}$. In practice, without directly manipulating, $\alpha_{i^{\prime},t^{\prime}}>0$ holds for many noisy messages $\mathbf{v}_{i^{\prime},t^{\prime}}$, as shown in Figure 5(c)-5(d) in our experiments.
@@
Let $\mathcal{Q}=\mathbf{h}_{i,t}^{in}$ be the target query embedding, and $\mathcal{N}(i)$ be the neighbors of the $i$-th sensor, i.e., $\mathbf{A}[i,i^{\prime}]\neq 0$, $\forall i^{\prime}\in\mathcal{N}(i)$. Denote the context keys as $\mathcal{K}=\{\mathbf{h}_{i^{\prime},t^{\prime}}^{in}\}_{i^{\prime},t^{\prime}}$ and the message values from $\mathcal{N}(i)$ to the point $(i,t)$ as $\mathcal{V}=\{\mathbf{v}_{i^{\prime},t^{\prime}}\}_{i^{\prime},t^{\prime}}$, where $i^{\prime}\in\mathcal{N}(i)$, $t^{\prime}\in\{1,\cdots,T\}$ and $\mathbf{v}_{i^{\prime},t^{\prime}}=\text{MLP}([\mathbf{h}_{i,t}^{in};\mathbf{h}_{i^{\prime},t^{\prime}}^{in}])$. We define SCA as: